# Polynomial time sampling from log-smooth distributions in fixed dimension under semi-log-concavity of the forward diffusion with application to strongly dissipative distributions

### v2

@@ Abstract
In this article, we provide a stochastic sampling algorithm with polynomial complexity in fixed dimension that leverages the recent advances on diffusion models where it is shown that under mild conditions, sampling can be achieved via an accurate estimation of intermediate scores across the marginals $(p_{t})_{t\geq 0}$ of the standard Ornstein-Uhlenbeck process started at $\mu$, the density we wish to sample from. The heart of our method consists into approaching these scores via a computationally cheap estimator and relating the variance of this estimator to the smoothness properties of the forward process. Under the assumption that the density to sample from is $L$-log-smooth and that the forward process is semi-log-concave: $-\nabla^{2}\log(p_{t})\succeq-\beta I_{d}$ for some $\beta\geq 0$, we prove that our algorithm achieves an expected $\epsilon$ error in KL divergence in $O(d^{7}(L+\beta)^{2}L^{d+2}\epsilon^{-2(d+3)}(d+m_{2}(\mu))^{2(d+1)})$ time with $m_{2}(\mu)$ the second order moment of $\mu$. In particular, our result allows to fully transfer the problem of sampling from a log-smooth distribution into a regularity estimate problem. As an application, we derive an exponential complexity improvement for the problem of sampling from an $L$-log-smooth distribution that is $\alpha$-strongly log-concave outside some ball of radius $R$: after proving that such distributions verify the semi-log-concavity assumption, a result which might be of independent interest, we recover a $poly(R,L,\alpha^{-1},\epsilon^{-1})$ complexity in fixed dimension which exponentially improves upon the previously known $poly(e^{LR^{2}},L,\alpha^{-1},\log(\epsilon^{-1}))$ complexity in the low precision regime.
@@ 1 Introduction
Given some potential $V:\mathbb{R}^{d}\to\mathbb{R}$ such that $\int e^{-V(x)}\mathrm{d}x<\infty$, the sampling problem consists into obtaining a sample from some distribution $p$ such that $p$ is $\epsilon$-close to $\mu\propto e^{-V}$ with respect to some divergence while maintaining the complexity, a.k.a. the number of queries to $V$ and possibly to its derivatives, as low as possible with respect to $\epsilon$ and the other parameters of the problem. As in Euclidean optimization, appealing complexity bounds can be obtained when $\mu$ is assumed to be $L$-log-smooth and $\alpha$-strongly log-concave or equivalently, when the potential $V$ is assumed to have an $L$-Lispchitz gradient and to be $\alpha$-strongly convex. Namely, denoting the condition number $\kappa=L/\alpha$, it has been recently shown that the Underdamped Langevin Algorithm (ULA) [14, 49], one of the most popular sampling algorithm, could provide a sample $X\sim p$ such that $\mathop{\mathrm{KL}}\nolimits(p,\mu)\lesssim\epsilon$ in $\tilde{O}(d^{1/3}\kappa\epsilon^{-2/3})$ time under minor modifications [43, 20, 7]. Similarly, it was shown that the Metropolis Adjusted Langevin Algorithm (MALA) [48, 3], a debiased version of Langevin, could achieve a complexity of $\tilde{O}(d^{1/2}\kappa\log^{4}(1/\epsilon))$ for a well-chosen warm start [2]; in both cases, the dependence is sub-linear in the dimension. However, because strong log-concavity implies unimodality, this framework is far from satisfactory when it comes to obtaining guarantees for real-world applications, where the distributions to sample from can be highly multimodal [30, 17]. One natural relaxation is to confine mutlimodality within some fixed ball, also known as the strong dissipativity assumption [13, 31, 32, 19] where we assume that the distribution is strongly log-concave outside this ball. Note that this assumption is much more realistic as it encompasses several types of Gaussian mixtures and other popular Bayesian mixture models [34, 33, 19]. Yet, this extra flexibility currently comes at a heavy price as known complexity bounds now degrade exponentially with the size of the non-convex region: if we assume that $\mu$ is $\alpha$-strongly log-concave outside the centered ball of radius $R$, ULA achieves a complexity of $\tilde{O}(e^{32LR^{2}}\kappa^{2}\frac{d}{\epsilon^{2}})$ and MALA achieves a complexity of $O(\frac{e^{40LR^{2}}}{\alpha}\kappa^{3/2}d^{1/2}(d\log(\kappa)+\log(1/\epsilon))^{3/2})$ [31]. Note that even if those are indeed worst case bounds, it has been well observed in practice that the aforementioned samplers suffer from metastability: as the distance between the modes grows, samples get stuck in one local mode, leaving the other ones unexplored [44, 26] (see also Chehab et al., [9] for formal lower bounds in continuous time). While there is no hope for a general polynomial bound as it was shown in Lee et al., [29] that even sampling from a $2$-modes distribution could require up to $2^{cd}$ queries to $V$ with $c$ a universal constant, sampling from a mutlimodal density in low dimension with provable polynomial bounds is still an open problem. This raises the following informal question:
“In fixed dimension, can we sample from a multimodal distribution in polynomial time?”
1.1 Related work
Sampling from multimodal distributions has been an active area of research for several decades now [37, 15, 38]. Indeed, it has long been observed that Monte Carlo Markov Chain (MCMC) based on local exploration of the space, such as ULA and MALA, are prawn to metastability and remain stuck in local modes [47, 21, 42]. We review in this paragraph the main alternatives to Langevin based samplers that have been developed and the guarantees they provide.
Histogram methods
Given some compact set $S\subset\mathbb{R}^{d}$, histogram methods build a regular $n$-grid over $S$ and sample a point $s_{i}$ of the grid according to the probability $e^{-s_{i}}/\sum_{i=1}^{n}e^{-V(s_{j})}$. If $V$ is assumed to be $\alpha$-strongly convex and $L$-smooth for instance, then, if $S$ is taken to be a ball of radius $r$ centered at $x^{*}$ the global minimum of $V$, the typical error of this method scales in $O(L(r/n)^{1/d}r+e^{-\alpha r^{2}})\,,$ thus inducing a typical complexity of $\tilde{O}((L/\epsilon)^{d})$ which is indeed polynomial in fixed dimension. However, if instead of being centered at the global minimum $x^{*}$, the ball is centered at some point $x$, the error remains $O(1)$ as long as $r\leq\|x-x^{*}\|$ as only the tail of $\mu$ is now discretized. When $V$ is smooth and strongly convex, $x^{*}$ can be approximated up to arbitrary precision yet in the case where $V$ is assumed to be strongly convex only outside of some unknown ball, one cannot localize the global minimum $x^{*}$ even in one dimension. In other terms, unless we have prior knowledge on the localization of the mass of $\mu$, histogram methods can induce an arbitrarily large error.
Importance Sampling (IS)
Instead of focusing on the problem of sampling from $\mu$, IS focuses on the related, yet slightly simpler problem of approximation of integrals under $\mu$: given some bounded function $f$, IS methods approximate the integral $\mu(f)\coloneqq\int_{x}f(x)\mathrm{d}\mu(x)\,.$ For any $W$ such that $\int e^{-W(x)}\mathrm{d}x<+\infty$, denoting $\nu\propto e^{-W}$, the integral above can be re-written as an expectation under $\nu$ as $\mu(f)=\mathbb{E}_{Y\sim\nu}[f(Y)\frac{\mathrm{d}\mu}{\mathrm{d}\nu}(Y)]\,.$ In particular, if we can access $n$ i.i.d. samples $(x_{i})_{1\leq i\leq n}$ from the proposal $\nu$, this quantity can be approximated as $\hat{\mu}(f)=\frac{\sum_{i=1}^{n}f(x_{i})e^{-V(x_{i})+W(x_{i})}}{\sum_{i=1}^{n}e^{-V(x_{i})+W(x_{i})}}\,.$ It was shown in Robins et al., [41] that the average quadratic error of this estimator scales as $O\biggl{(}\frac{\chi^{2}(\mu,\nu)}{n}\biggr{)}\,.$ Yet, unless $\nu$ is carefully chosen, this bound may blow up to infinity even for simple densities: if $\mu\sim\mathcal{N}(\mu_{1},\sigma_{1}^{2})$ and $\nu\sim\mathcal{N}(\mu_{2},\sigma_{2}^{2})$, it holds that $\chi^{2}(\mu,\nu)=\frac{1}{2}\biggl{(}\frac{\sigma_{2}^{2}}{\sigma_{1}\sqrt{2\sigma_{2}^{2}-\sigma_{1}^{2}}}e^{\frac{(\mu_{1}-\mu_{2})^{2}}{2\sigma_{2}^{2}-\sigma_{1}^{2}}}-1\biggr{)}\,.$ In particular, not only this bound blows up when the variance of $\nu$ is not at least half the variance of $\mu$ but also, this bound always degrades exponentially with the quadratic distance between the mode of the proposal and the mode of the target $\mu$. Hence, even in this one dimensional strongly log-concave toy example, IS requires a precise a priori knowledge of $\mu$ to tailor a good proposal $\nu$ (and also requires to be able to sample from $\nu$). Nevertheless, we discuss in the last paragraph of this section the recent work of Chak, [8] who shows that if the localization of the mass of $\mu$ is known, as in histogram methods, a suitable proposal $\nu$ can be designed with appealing bounds.
Tempering heuristics
@@ 1.1 Related work
Instead of directly sampling from $\mu\propto e^{-V}$, we start to sample from a flattened version of $\mu$ given by $\mu^{\beta}\propto e^{-\beta V}$ for $\beta$ small and we progressively increase $\beta$ to $1$. When $\mu$ is assumed to be a finite mixture of the same shifted $\alpha$-strongly log-concave and $L$-log-smooth distribution, Lee et al., [29] proved that for a well-chosen (stochastic) sequence of flattened distributions $\mu^{\beta_{i}}$, sampling up to precision $\epsilon$ in $\mathop{\mathrm{KL}}\nolimits$ can be achieved in $poly(L,\alpha^{-1},w_{min},d,\epsilon^{-1},R)$ time with $R$ the diameter of the shifts and $w_{\min}$ the minimum weight in the mixture. Yet, despite the strong guarantees it offers, this framework excludes many practical cases such as mixtures of Gaussians with different variances for instance: in that case, the tempering strategy even fails in low dimension as it attributes biased weights to the different modes.
Reverse Diffusion (RD) methods
As we shall review in detail in Section 2, recent works show that, under milder and milder assumptions [12, 10, 16, 5], the problem of sampling can be transferred into the problem of approximating the score in $L^{2}$ distance along the standard Ornstein-Uhlenbeck (OU) process initialized at the targeted distribution $\mu$. Over the past year, several works sought to leverage this result and focused on the task of estimating these intermediate scores given an oracle access to $V$ and eventually to its gradient. Their guarantees can be summarized as follows: in Ding et al., [18], denoting $\pi^{\eta,\Sigma}$ the $d$-dimensional Gaussian density with mean $\eta$ and covariance matrix $\Sigma$, the authors proved that if there existed known $\eta,\Sigma$ such that $\mathrm{d}\mu/\mathrm{d}\pi^{\eta,\Sigma}$ is globally $L$-Lipschitz and strictly bounded from below by $\xi$, then their algorithm achieved a complexity $poly(L,\xi^{-1},d,\epsilon^{-1})$. We show in Appendix C that even for the simple mixture $\mu=\frac{1}{2}\mathcal{N}(1,1)+\frac{1}{2}\mathcal{N}(-1,1)\,,$ the previous conditions do not hold. In Huang et al., [25], it is shown that under the assumption that the OU process remains $L$-log-smooth along the trajectory, denoting $m_{2}(\mu)=\int\|x\|^{2}\mathrm{d}\mu(x)$ the second moment of $\mu$, their algorithm could reach a complexity of $e^{O(L^{3}\log^{3}((Ld+m_{2}(\mu))/\epsilon))}$. In He et al., [22], it is shown that under a sub-quadratic growth assumption on $V$ (which is weaker than log-smoothness), i.e. there exists $L>0$ such that $V(x)-V(x^{*})\leq L\|x-x^{*}\|^{2}\,,$ with $x^{*}$ a global minimizer of $V$, their algorithm could reach at best a complexity of $O(L^{d/2}\epsilon^{-d}e^{(L\|x^{*}\|^{2}+\|x_{N}\|^{2})})\,,$ where $x_{N}$ is the sample output by their algorithm. Under the reasonable (and desirable) assumption that $\mathbb{E}[||x_{N}||^{2}]\asymp m_{2}(\mu)$, Jensen inequality yields that their complexity scales at best as $O(L^{d/2}\epsilon^{-d}e^{(L\|x^{*}\|^{2}+m_{2}(\mu))})\,.$ Hence, we observe that previous approaches based on reverse diffusion suffer at best from exponential complexity either with respect to $L$ or to $m_{2}(\mu)$.
1.2 Our contributions
Polynomial time sampling under semi-log-concavity of the Ornstein-Uhlenbeck (OU) process
As in the aforementioned work, we rely on the recent advances on reverse diffusion and focus on the task of estimating the intermediate scores $\nabla\log(p_{t})$ with $(p_{t})_{t\geq 0}$ the marginals of the standard OU process initialized at $\mu$. Yet, unlike all previous works, we provide fully polynomial time guarantees in fixed dimension when $\mu$ is assumed to be log-smooth and when the forward process has semi-log-concave marginals.
Theorem 1 (Main result, informal)

### v3

# Sampling from multi-modal distributions with polynomial query complexity in fixed dimension via reverse diffusion

@@ Abstract
Even in low dimensions, sampling from multi-modal distributions is challenging. We provide the first sampling algorithm for a broad class of distributions — including all Gaussian mixtures — with a query complexity that is polynomial in the parameters governing multi-modality, assuming fixed dimension. Our sampling algorithm simulates a time-reversed diffusion process, using a self-normalized Monte Carlo estimator of the intermediate score functions. Unlike previous works, it avoids metastability, requires no prior knowledge of the mode locations, and relaxes the well-known log-smoothness assumption which excluded general Gaussian mixtures so far.
@@ 1 Introduction
Sampling from a distribution whose density is only known up to a normalization constant is a fundamental problem in statistics. Formally, given some potential $V:\mathbb{R}^{d}\to\mathbb{R}$ such that $\int e^{-V(x)}\mathrm{d}x<\infty$, the sampling problem consists in obtaining a sample from some distribution $p$ such that $p$ is $\epsilon$-close to the target $\mu\propto e^{-V}$ with respect to some divergence while maintaining the complexity, i.e., the number of queries to $V$ and possibly to its derivatives, as low as possible. Depending on the shape of the distribution, the typical complexity of existing sampling algorithms can significantly differ.
Log-concave and "PL-like" distributions
As in Euclidean optimization, a common assumption in the sampling literature is to assume that $\mu$ is log-concave and log-smooth or equivalently, that $V$ is convex and smooth. Specifically, when the potential $V$ is assumed to be $\alpha$-strongly convex and to have an $L$-Lipschitz gradient, the popular Unajusted Langevin Algorithm (ULA) is known to achieve fast convergence [Durmus and Moulines, 2017, Dalalyan and Karagulyan, 2019]. Because strong-log concavity implies uni-modality, thus excluding many distributions of interest, this assumption was further relaxed to $\mu$ verifying an ${\alpha}^{-1}$-log-Sobolev inequality, later interpreted as a Polyak–Łojasiewicz type condition on $\text{KL}(\cdot|\mu)$ for the Wasserstein geometry [Blanchet and Bolte, 2018]. Under these conditions, ULA was shown to achieve $\epsilon$-error in Kullback-Leibler ($\mathop{\mathrm{KL}}\nolimits$) divergence in $\tilde{O}(L^{2}\alpha^{-2}d\epsilon^{-1})$ queries to $\nabla V$ [Vempala and Wibisono, 2019]. While these polynomial guarantees do go beyond the uni-modal setting, we show in the next paragraph that most existing algorithms still fail to sample from truly multi-modal distributions.
Multi-modal distributions
Designing sampling algorithms for multi-modal distributions is an active area of research. However, most existing sampling algorithms are limited in at least two of the following ways:
  1. 1.
    The query complexity is exponential in the parameters of the problem. For instance, when the global $\alpha$-strong convexity of $V$ is relaxed with $\alpha$-strong convexity outside a ball of radius $R$, the log-Sobolev constant of $\mu$ degrades to $O(e^{8LR^{2}}\alpha)$ [Ma et al., 2019, Prop. 2]. In practice, this does not simply translate to poor worst case bounds: practitioners are well aware that when dealing with multi-modal distributions, ULA-based algorithms suffer from metastability, where they get stuck in local modes, leading to slow convergence [Deng et al., 2020].
  2. 2.
    Guarantees are obtained under a log-smoothness assumption; we show in Sec. 3.2 that even for Gaussian Mixtures, which is arguably the most basic multi-modal model, this assumption may not be verified.
  3. 3.
    Explicit a priori knowledge on the target distribution is required for the algorithm to converge. For instance, importance sampling requires a proposal distribution whose "effective" support must cover sufficiently well the one of the target; however, since $V$ can only be evaluated point-wise, the access to such a support is unclear. In fact, this limitation goes beyond the multi-modal setting in some cases: no matter the log-Sobolev constant of the target, similarly to euclidean gradient descent, ULA still requires an a priori bound on the inverse smoothness constant of $V$ and sufficiently small step size in comparison in order to converge.
Generally speaking, finding a sampling algorithm that addresses even the first limitation is not possible. Recent results provide lower-bounds on the complexity of sampling from a multi-modal distribution: they are exponential in the dimension, as shown in Lee et al. [2018, Th K.1], Chak [2024, Th 3.3] and He and Zhang [2025, Th. 3]. However, when the dimensionality is fixed, the question of whether the three limitations can be addressed remains. In fixed dimension, can we sample from a broad class of multi-modal distributions with a polynomial number of queries in the parameters of the problem, (e.g. $L,R,\alpha$) and without prior knowledge on these parameters?
Our contributions
Our work answers this question positively. We provide a sampling algorithm that addresses the three limitations outlined above. First, we show that our algorithm has polynomial complexity in all the problem parameters but the dimension, thus enabling to efficiently sample from highly multi-modal distributions in fixed dimension. Second, we show that unlike most existing results, our guarantees hold under relaxed regularity assumptions that cover general Gaussian mixtures. Third, our algorithm yield guarantees without prior knowledge on the parameters of the distribution.
The workhorse of our algorithm is made of two key ingredients: the reverse diffusion scheme, that transfers the sampling problem into a score estimation problem with different levels of noise, and self-normalized importance sampling, to estimate these noisy scores using only query access to $V$ and additional Gaussian samples. Both reverse diffusion [Huang et al., 2024a, b, He et al., 2024] and self-normalized importance sampling [Huang et al., 2025, Jiao et al., 2021, Ruzayqat et al., 2023, Ding et al., 2023, Saremi et al., 2024] were already used separately to sample in a Bayesian setting, where only the log-density of the target is available. However, these works either offer no theoretical guarantees or fail to improve upon existing results. In this article, we combine both reverse diffusion and self-normalized importance sampling in a single algorithm, allowing us to recover the first polynomial time guarantees to sample from highly multi-modal distributions in fixed dimension.
This paper is structured as follows. First, we survey related work on multi-modal sampling in Section˜2 and present our main result in Section˜3. Then, we detail our sampling algorithm in Section˜4 and provide the key ingredients for the proof in Section˜5.
2 Related work
Before detailing our sampling algorithm in Section 4, we review the main alternatives to ULA when sampling from a multi-modal distribution and the guarantees they offer. We show that existing approaches suffer from at least two of three drawbacks outlined in the introduction: exponential query complexity, a restrictive smoothness assumption that excludes general Gaussian mixtures (as shown later in Sec. 3.2), and unavailable prior knowledge of the distribution.
Proposal-based algorithms
At a high level, proposal-based algorithms use a proposal distribution that is easy to sample from and that can either be directly used as a proxy for the target, or, alternatively, whose samples will be rejected (respectively re-weighted) to obtain approximate samples from the target; the corresponding algorithm for the latter case is the well-known rejection sampling (respectively importance sampling) scheme. Guarantees for these methods typically assume that the target distribution is log-smooth. Furthermore, they must use carefully designed proposals that require prior knowledge of the target distribution that is often unavailable in practice.
For instance, when the target is $L$-log-smooth and with finite second moment $m_{2}$, one can design and sample from a proxy distribution that is $\epsilon$-close in $\mathop{\mathrm{TV}}\nolimits$ to the target using $\tilde{O}((Lm_{2}\epsilon^{-1})^{O(d)})$ queries to the potential function $V$ and its gradient $\nabla V$ [He and Zhang, 2025]. While this bound is indeed polynomial in fixed dimension, the design of the proxy requires an $\epsilon$-approximation of the global minimum of $V$ which can only be achieved if $m_{2}$, therefore the location of the mass, is explicitly available; this is rarely the case in practice.
Similarly, if we further assume that the target is $\alpha$-strongly log-concave outside of a ball of radius $R$, one can achieve an $\epsilon$-precise approximation of the target distribution in $\mathop{\mathrm{TV}}\nolimits$ with polynomial number of queries to $\nabla V$ when $d$ is fixed via an importance sampling scheme [Chak, 2024, Th. 2.3]. In this case, the proposal is such that it coincides with the target when the potential of the target is above some cutoff, and is flat elsewhere. While there exists a cutoff value ensuring the log-concavity of the proposal, thus allowing its efficient sampling, this value depends on the unknown constants $L,R$ and $\alpha$ [Chak, 2024, Prop. 5.1].
Tempering-based algorithms
@@ 2 Related work
Instead of directly sampling from $\mu\propto e^{-V}$, these algorithms start by sampling from a flattened version of $\mu$ given by $\mu^{\beta}\propto e^{-\beta V}$ for small $\beta$, and gradually increase $\beta$ to $1$, a strategy sometimes also referred to as annealing. When $\mu$ is assumed to be a finite mixture of the same shifted $\alpha$-strongly log-concave and $L$-log-smooth distribution, Lee et al. [2018] proved that for a well-chosen (stochastic) sequence of flattened distributions $\mu^{\beta_{i}}$, sampling up to precision $\epsilon$ in $\mathop{\mathrm{KL}}\nolimits$ can be achieved in $poly(L,\alpha^{-1},w_{min}^{-1},d,\epsilon^{-1},R)$ queries to $\nabla V$, where $R$ is the location of the furthest mode and $w_{\min}$ the minimum weight in the mixture. However, this setting is quite restrictive: for Gaussian mixtures for instance, it only handles the case where the covariance matrices are identical for all components. Furthermore, the algorithm requires explicit knowledge of $R$ which is unavailable.
Föllmer Flows
Instead of directly sampling from the target distribution $\mu\propto e^{-V}$, Föllmer Flows start from a simple (e.g. Gaussian or Dirac mass) distribution that is progressively interpolated to the target using a Schrödinger bridge. When the initial distribution is a Dirac mass at 0, this bridge solves a closed-from SDE [Wang et al., 2021, Theorem 3] which can thus be discretized to generate samples from the target. In the context of Bayesian inference, where one only has access to the unnormalized density, Vargas et al. [2022] estimated the drift of the resulting SDE via neural methods; in particular no guarantees on the sampling quality are provided. In Huang et al. [2025], Jiao et al. [2021], Ruzayqat et al. [2023], Ding et al. [2023], the drift is estimated via a Monte-Carlo method. While in appearance, these works provide strong polynomial guarantees, a closer look shows that these guarantees only hold if the function $f(x)=e^{-V(x)+\|x\|^{2}/2}$ is Lipschitz, smooth and bounded from below; we show in Appendix B.1.2 that this assumption is quite restrictive.
Diffusion-based methods
Over the past few years, diffusion-based algorithms, and especially reverse diffusion, that we shall review in details in Sec. 3, have emerged as solid candidates for multi-modal sampling. In essence, they allow to transfer the sampling problem into the problem of estimating the scores of intermediate distributions that are given by the convolution of the initial distribution with increasing levels of Gaussian noise. Under an $\epsilon$-oracle of these intermediate scores, it has been shown that diffusion-based methods could yield an $\epsilon$-approximate sample of the target in $\text{poly}(\epsilon^{-1})$ time under milder and milder assumptions [Chen et al., 2023b, a, Benton et al., 2024, Li et al., 2024, Conforti et al., 2025, Gentiloni-Silveri and Ocello, 2025, Cordero-Encinar et al., 2025]. This framework has been applied with tremendous empirical success in generative modeling, where numerous samples of the target are already available and one seeks to produce new samples from the target, for several years now [Song and Ermon, 2019, Ho et al., 2020, Bortoli et al., 2021, Chen et al., 2024]. However, it was only quite recently that this framework has been applied in a Bayesian context, where only an unnormalized density of the target, instead of samples, is available [Huang et al., 2024a, b, He et al., 2024, Grenioux et al., 2024, Akhound-Sadegh et al., 2024]. In a work closely related to ours, Huang et al. [2024b] showed that if the intermediate scores remain $L$-log-smooth for any noise level, which in particular implies that the target itself is $L$-log-smooth, their algorithm could reach a complexity of $O(e^{L^{3}\log^{3}((Ld+m_{2})/\epsilon)})$ with $m_{2}$ the second order moment of the target. In He et al. [2024], the smoothness assumption is relaxed with a sub-quadratic growth assumption $V(x)-V(x^{*})\leq L\|x-x^{*}\|^{2}$, where $x^{*}$ is any global minimizer of $V$, yet at the price of an oracle access to $V(x^{*})$; under these assumptions, the authors manage to obtain a complexity that is at best $O(L^{d/2}\epsilon^{-d}e^{L\|x^{*}\|^{2}+\|x_{N}\|^{2}})$ where $x_{N}$ is the final output sample. Under the reasonable (and desirable) assumption that $\mathbb{E}[\|x_{N}||^{2}]\approx m_{2}$, Jensen’s inequality yields an overall $O(L^{d/2}\epsilon^{-d}e^{L\|x^{*}\|^{2}+m_{2}})$ complexity. In particular, both these works suffer at least two of the limitations mentioned in the introduction, making them ill-suited for multi-modal sampling.
3 Presentation of the main result and application to Gaussian Mixtures
3.1 Main result
As in the works mentioned above, we rely on the recent advances on reverse diffusion and focus on the task of estimating the intermediate scores. Using an estimator that is described in Sec. 4, we recover polynomial sampling guarantees for densities verifying the assumptions described hereafter.
Assumption 1 (Semi-log-convexity)
We assume that $\mu\propto e^{-V}$ is such that $\log(\mu)$ is $C^{2}$ and verifies $\nabla^{2}\log(\mu)\succeq-\beta I_{d}$ or equivalently $\nabla^{2}V\preceq\beta I_{d}$ for some $\beta\geq 0$.
This assumption shall be referred to as semi-log-convexity, by analogy with the functional analysis literature [Mikulincer and Shenfeld, 2023, Theorem 3]. Note that is has also been referred to as 1-sided Lipschitzness for $\nabla V$ [Gentiloni-Silveri and Ocello, 2025]. It relaxes the classical log-smoothness assumption which implies the additional lower bound $\nabla^{2}V\succeq-\beta I_{d}$. In particular, unlike the latter, a mixture of semi-log-convex densities remains semi-log-convex [Marshall et al., 1979, Chap. 16.B]; we provide a quantitative version of this statement in Sec. 3.2 in the case of Gaussian mixtures.
Assumption 2 (Dissipativity)
We assume that $\mu\propto e^{-V}$ is such that there exists $a>0,b\geq 0$ for which its potential satisfies $\langle\nabla V(x),x\rangle\geq a\|x\|^{2}-b$ for all $x\in\mathbb{R}^{d}$.
This assumption is referred to as dissipativity as common in the sampling and optimization literature [Raginsky et al., 2017, Zhang et al., 2017, Erdogdu and Hosseinzadeh, 2021]. Note that this assumption relaxes strong convexity outside of a ball which can be equivalently re-written as $\langle\nabla V(x)-\nabla V(y),x-y\rangle\geq a\|x-y\|^{2}-b$ for all pairs $(x,y)\in\mathbb{R}^{d}$, also referred to as strong dissipativity [Eberle, 2013, Erdogdu et al., 2022]. We show in Sec. 3.2 that unlike strong-dissipativity, a mixture of dissipative distributions remain dissipative. distribution is also log-concave outside a ball.
Theorem 1